# THE POLARIZATION AND MAGNETIC FIELD OF THE RADIO ARC AS OBSERVED BY ALMA AT 100 GHZ

### v1

@@ V Conclusions
In this work we have presented the first ever polarimetric ALMA observations of the unique GC NTFs which targeted the prominent Radio Arc NTF. This also is the first time sub-arcsecond polarimetric observations have been conducted for the NTFs at 100 GHz. We derived the polarized intensity and magnetic field distributions for the Radio Arc from these observations, finding a magnetic field that is generally rotated with respect to the orientation of the Radio Arc NTFs but is uniform along the lengths of the Radio Arc filaments. These results build on previous VLA results showing an alternating magnetic field pattern for the Radio Arc where the field alternates between being parallel and rotated with respect to the Radio Arc orientation. We summarize the major findings of this work here:
The total intensity of the Radio Arc from ALMA (shown in the upper panel of Figure 1) accurately depicts the Radio Arc filaments based on previous radio observations conducted with instruments like the VLA (e.g., YMC1984; Pare2019; Heywood2022).
We find significant polarized intensity in the southern portion of the Radio Arc, with some of the polarization lacking a significant total intensity counterpart as shown in Figure 2 (labeled as PF2 in that figure). The features of this polarized intensity distribution are similar to previous radio polarimetric observations of the Radio Arc that reveal extensions of polarizations lacking total intensity filament counterparts (Inoue1989; Pare2019). These features are likely caused by a foreground structure that is coherent in total intensity on scales larger than the maximum recoverable scale of our interferometric observations.
The magnetic field inferred from our ALMA observations lacks the alternating pattern observed from previous 10 GHz VLA observations (Pare2019). However, the field inferred from ALMA is ubiquitously rotated with respect to the Radio Arc filament orientation. The field orientation is also systematically different between PF1 and PF2 as shown in Figure 3.
We update our hypothesis for the line-of-sight geometry of structures local to the Radio Arc based on the ALMA magnetic field results presented here, as shown in panels b and c of Figure 4, building on the sketch presented in Pare2019. We propose that sketch c is the most accurate depiction of reality, where the 100 GHz ALMA observations are most effected by the central concentration of polarization within the NTF filaments but are less sensitive to the Radio Shell density enhancements. This explains the different magnetic field distributions inferred for the VLA and ALMA observations.

### v2

@@ V Conclusions
In this work we have presented the first ever polarimetric ALMA observations of the unique GC NTFs which targeted the prominent Radio Arc. This also is the first time sub-arcsecond polarimetric observations have been conducted for the NTFs at 100 GHz. We derived the polarized intensity and magnetic field distributions for the Radio Arc from these observations, finding a magnetic field that is generally rotated with respect to the orientation of the Radio Arc NTFs but is uniform along the lengths of the Radio Arc filaments. These results build on previous VLA results showing an alternating magnetic field pattern for the Radio Arc where the field alternates between being parallel and rotated with respect to the Radio Arc orientation. We summarize the major findings of this work here:
We find significant polarized intensity in the southern portion of the Radio Arc, with some of the polarization lacking a significant total intensity counterpart as shown in Figure 2 (labeled as PF2 in that figure). The features of this polarized intensity distribution are similar to previous radio polarimetric observations of the Radio Arc that reveal extensions of polarizations lacking total intensity filament counterparts (Inoue et al., 1989; Paré et al., 2019). These features are likely caused by a foreground structure that is coherent in total intensity on scales larger than the maximum recoverable scale of our interferometric observations.
The magnetic field inferred from our ALMA observations lacks the alternating pattern observed from previous 10 GHz VLA observations (Paré et al., 2019). However, the field inferred from ALMA is ubiquitously rotated with respect to the Radio Arc filament orientation. The field orientation is also systematically different between PF1 and PF2 as shown in Figure 3.
We update our hypothesis for the line-of-sight geometry of structures local to the Radio Arc based on the ALMA magnetic field results presented here, as shown in panels b and c of Figure 4, building on the sketch presented in Paré et al. (2019). We propose that sketch c is the most accurate depiction of reality, where the 100 GHz ALMA observations are most affected by the central concentration of polarization within the NTF filaments but are less sensitive to the Radio Shell density enhancements. This explains the different manifestations of the magnetic field distribution inferred for the VLA and ALMA observations.